Lou Barlow, a divisional manager for Sage Company, has an opportunity to manufacture and sell one of two new products for a five-year period. His annual pay raises are determined by his division’s return on investment (ROI), which has exceeded 22% each of the last three years. He has computed the cost and revenue estimates for each product as follows:

Product A Product B
Initial investment:
Cost of equipment (zero salvage value) $ 370,000 $ 570,000
Annual revenues and costs:
Sales revenues $ 400,000 $ 480,000
Variable expenses $ 180,000 $ 214,000
Depreciation expense $ 74,000 $ 114,000
Fixed out-of-pocket operating costs $ 88,000 $ 68,000

The company’s discount rate is 20%.

Click here to view Exhibit 13B-1 and Exhibit 13B-2, to determine the appropriate discount factor using tables.

Required:

1. Calculate the payback period for each product.

2. Calculate the net present value for each product.

3. Calculate the internal rate of return for each product.

4. Calculate the project profitability index for each product.

5. Calculate the simple rate of return for each product.

6a. For each measure, identify whether Product A or Product B is preferred.

1) Calculate the payback period for each product. (Round your answers to 2 decimal places.

Product A Product B
Payback period years    years

2) Calculate the net present value for each product. (Round your final answers to the nearest whole dollar amount.)

Product A Product B
Net present value

3)Calculate the internal rate of return for each product. (Round your answers to 1 decimal place i.e. 0.123 should be considered as 12.3%.

Product A Product B
Internal rate of return % %

4) Calculate the project profitability index for each product. (Round your answers to 2 decimal places.

Product A Product B
Project profitability index

5) Calculate the simple rate of return for each product. (Round your answers to 1 decimal place i.e. 0.123 should be considered as 12.3%.)

Product A Product B
Simple rate of return    %    %

6A) For each measure, identify whether Product A or Product B is preferred.

Net Present Value Profitability Index Payback Period Internal Rate of Return Simple Rate of Return
Answer #1
Product A Product B
1 Payback period =                      2.80           2.88
2 Net Present Value = $             24,812 $ 22,218
3 IRR = 23.01% 21.75%
4 Profitability Index                      1.07           1.04
5 Simple rate of return = 31.4% 29.5%
6a Net Present Value = Product A
Profitability Index = Product A
Payback period = Product A
IRR = Product A
Simple rate of return = Product A
Workings:
Product A:
2
Year Value Flows Present Factor @20% Present Value
Initial Cost 0 $       -3,70,000 1 $       -3,70,000
Cash Inflows ($400000-$180000-$88000) 1 - 5 $         1,32,000 2.991 $         3,94,812
Net Present Value $            24,812
Product B:
Year Value Flows Present Factor @20% Present Value
Initial Cost 0 $       -5,70,000 1 $       -5,70,000
Cash Inflows ($480000-$214000-$68000) 1 - 5 $         1,98,000 2.991 $         5,92,218
Net Present Value $            22,218
1 Product A:
Payback period = Initial investment / Annual net cash flows
= $370000 / $132000
=                      2.80 years
Product B:
Payback period = Initial investment / Annual net cash flows
= $570000 / $198000
=                      2.88 years
3 Computation of IRR
Product A:
Year Value Flows
0 $       -3,70,000
1 $         1,32,000
2 $         1,32,000
3 $         1,32,000
4 $         1,32,000
5 $         1,32,000
IRR = 23.01%
Formula =IRR(I15:I20)
Product B:
Year Value Flows
0 $       -5,70,000
1 $         1,98,000
2 $         1,98,000
3 $         1,98,000
4 $         1,98,000
5 $         1,98,000
IRR = 21.75%
Formula =IRR(I15:I20)
4 Product A:
Profitability Index = 1 + NPV/Initial Investment
= 1 + $24812 /$370000
=                      1.07
Product B:
Profitability Index = 1 + NPV/Initial Investment
= 1 + ($22218 /$570000)
=                      1.04
5 Product A:
Computation of Simple rate of return:
Simple rate of return = Net Profit /Average Investment
= $58000 / $185000
= 31.4%
Net Profit = ($400000-$180000-$88000-$74000)
= $             58,000
Average Investment = ($370000+0)/2
= $         1,85,000
Product B:
Simple rate of return = Net Profit /Average Investment
= $84000 / $285000
= 29.5%
Net Profit = ($480000-$214000-$68000-$114000)
= $             84,000
Average Investment = ($570000+0)/2
= $         2,85,000
6a Net Present Value = Product A
Profitability Index = Product A
Payback period = Product A
IRR = Product A
Simple rate of return = Product A